OpenStudy (anonymous):

I Will Give A Medal! A company produces accessories for smart phones and tablets. The profit on each smart phone case is $2 and the profit on each tablet case is $3. The company made a profit of $1,200 on the cases last month. The equation 2x + 3y = 1,200 represents the company's profit from cases last month, where x is the number of smart phone cases sold and y is the number of tablet cases sold. 2.Describe how you would graph this line using the slope-intercept method. Be sure to write in complete sentences.

OpenStudy (anonymous):

use a bar graph to line them out

OpenStudy (anonymous):

i also need help with this question 3.Write the equation in function notation. Explain what the graph of the function represents. Be sure to use complete sentences.

OpenStudy (ranga):

2x + 3y = 1,200 The slope-intercept form of the equation is: y = mx + b where m is the slope and b is the y-intercept. So we need to solve for y. Subtract 2x from both sides: 3y = -2x + 1200 divide both sides by 3 y = -2/3x + 400 Slope m = -2/3 and y-intercept = 400 y-intercept means it is where the graph will cross the y-axis. So you got one point on the straight line at (0, 400) Here the slope is negative. That means as x increases, y decreases. slope = rise / run = -2/3 so of you take 3 steps along the x-axis and the y goes down by 2 units. So from the first point (0,400) on the y-axis, take three units to the right and drop two units down to get your second point. Using a ruler draw a line passing through the first and second points.

OpenStudy (anonymous):

@ranga could you help me with these as well 4.Graph the function using one of the following two options below. One the graph, make sure to label the intercepts. You may graph your equation by hand on a piece of paper and scan your work. You may graph your equation using graphic technology that can be found in the Course Information area. 5.Suppose in the next month, the total profit on smart phone cases and tablet cases is $1,500. The profit amounts are the same, $2 for smart phone case and $3 for the tablet case. In a paragraph of at least three sentences, explain how the graphs of the functions for the two months are the same and how they are different. Be sure to use complete sentences. 6.Below is a graph that represents the total profits for a third month. Write the equation of the line that represents this graph. Show your work or explain how you determined the equations.

OpenStudy (anonymous):

here is the graph to go along with question 6

OpenStudy (ranga):

answering another question. will be here when done.

OpenStudy (ranga):

#4 Graph y = -2/3x + 400

OpenStudy (ranga):

#5) For the next month, the equation will be 2x + 3y = 1500 or y = -2/3x + 500 

OpenStudy (ranga):

Both lines will have the same slope of -2/3. Th graphs of the two functions will be two parallel lines. In both cases, as x increases, y decreases and vice-versa. The difference will be in the y-intercepts and the x-intercepts. The first graph has a y-intercept of 400 and the second graph has a y-intercept of 500. The x-intercepts are 600 and 750 respectively.

OpenStudy (ranga):

#6) The equation of a straight line in intercept form is: x/a + y/b = 1 where a is the x-intercept and b is the y-intercept. From the graph we can see the y-intercept b = 300 and the x-intercept a = 450. x/450 + y/300 = 1 The least common multiple of 450 and 300 is 900. Multiply the equation throughout by 900: 2x + 3y = 900 is the equation of the line shown in the diagram for #6.